A uniform plank ,10m long is balances on a pivot 2m from one end by applying a force of 30N at that end. What is the weight of the plank
1
Expert's answer
2018-06-10T10:29:08-0400
At the equilibrium the sum of moments about pivot is zero:
W(10/2-2)-F(2)=0
W=(2/3)F=(2/3)30=20 N.